The no of 3 digit nos. having only two consecutive digits identical is?

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear jee king

identical digits can be at 100th and 10th position  or at 10th and unit place

case 1

  identical digit at 100th and 10th posiition

  0 can be at 100th position so these 2 position can be filled by 9 ways

and unit place can be filled by 9 ways

 total ways =9*9 =81

case 2

identicle digit at 10 th and unit position

 total ways =9*9 =81

 

so total ways =81+81 =162

 


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin